On January 1, 2013, Powell Company purchased a building and equipment that have the following useful lives, salvage value, and costs.

Building, 25-year estimated useful life, $4,000,000 cost, $400,000 salvage value

Equipment, 15-year estimated useful life, $600,000 cost, no salvage value

The building has been depreciated under the straight-line method through 2017. In 2018, Powell decided to change the total useful life of the building to 30 years. The equipment is depreciated using the straight-line method, but in 2018, the company decided to switch to the double-declining balance method of depreciation for the equipment and changed the salvage value to $60,000.

Instructions

(a)    Prepare the journal entry necessary to record the depreciation expense on the building in 2018.

(b)    Compute depreciation expense on the equipment for 2018.
Answer #1

Answer a.

Calculation of Depreciation on Building
Cost of building $4,000,000
Salvage Value $400,000
Depreciation value $3,600,000
Life of Building 25
Depreciation per year $144,000
Year Up to end of 2017 5
Accumulated Depreciation for 5 years = 5*144000 720000
Book value on 1st Jan 2018 = $4000000-$720000= $3,280,000
Revised total life of building 30
Remaining life of Building 25
Salvage value assumed to be same $400,000
Depreciation value $2,880,000
Depreciation per year = $2880000/25 $115,200
2018
Date Account Debit Credit
31-Dec Depreciation Expense $115,200
Accumulated Depreciation-Building $115,200
Depreciation Expense recorded for 2018

Answer b.

Calculation of Depreciation on Equipment
Cost of equipment $600,000
Salvage Value $0
Depreciation value $600,000
Life of Equipment 15
Depreciation per year $40,000
Year Upto end of 2017 5
Accumulated Depreciation for 5 years = 5*40000 200000
Book value on 1st Jan 2018 = $4000000-$720000= $400,000
Remaining life of Equipment 10
Salvage value assumed to be same $60,000
Rate of Depreciation unde DDB = 1*2*100/10 20%
Depreciation for 2018 = 20% of $400000 = $80,000
2018
Date Account Debit Credit
31-Dec Depreciation Expense $80,000
Accumulated Depreciation -Equipment $80,000
Depreciation Expense recorded for 2018

FEEL FREE TO ASK ANY CLARIFICATION IF ANY REQUIRED KINDLY PROVIDE FEED BACK BY THUMBS UP IF SATISFIED IT WILL BE HIGHLY APPRECIATED

Best efforts are made to provide best answer still some error may creep in if something such happens kindly let me know i ll correct the error.

THANK YOU